# Prompt and Parameter Co-Optimization for Large Language Models

### v1

@@ 5 Related Work
Fine-tuning. Fine-tuning is a standard approach for adapting pre-trained language models to specific tasks. Recent advances in parameter-efficient fine-tuning (PEFT) scale this to LLMs by training small plug-in modules under a frozen backbone. Representative methods include Adapter Tuning (ref:adapter_tuning, ), which inserts lightweight adapter layers, and LoRA (ref:lora, ), which applies low-rank updates to weight matrices. Building on these, recent works propose various post-training strategies (ref:kto, ; ref:online_dpo, ; ref:bco, ; ref:cpo, ; ref:opro, ). For instance, RLHF (ref:rlhf, ) incorporates human feedback to improve alignment, while DPO (ref:dpo, ) streamlines preference modeling through contrastive learning.
Prompt optimization. Prompt optimization has become a research hotspot with the rise of in-context learning. Early works focus on continuous prompt tuning (ref:prefix_tuning, ; ref:google_soft_prompt, ; ref:p_tuning, ; ref:p_tuning_v2, ; ref:prompt_tuning, ), where learnable vectors are attached to the input. To mitigate the interpretability challenge (ref:hard_better_1, ), recent studies turn to discrete prompt optimization, which generates prompts in natural language. Existing approaches fall into two categories. The first employs LLMs as prompt optimizers (ref:hmaw, ), typically in a feedback-driven manner, via strategies such as reflection-based refinement (ref:opro, ; ref:llm_dual_phase_po, ; ref:grad_sum, ) or Monte Carlo Tree Search (MCTS) (ref:prompt_agent, ; ref:protegi, ; ref:cfpo, ). The second (ref:qpo, ; ref:uniprompt, ; ref:rlprompt_improve, ; ref:prompt_agent, ; ref:mapo, ) trains an auxiliary model to generate better prompts, most often with reinforcement learning, since prompts lack explicit ground truth and can be naturally guided by the task rewards.
Joint optimization. Recently, BetterTogether (ref:better_together, ) proposes to update prompts and model weights together to improve the task performance. In BetterTogether, prompt optimization and fine-tuning correspond to discrete and continuous optimization problems, respectively. During the prompt optimization phase, it samples candidate demonstration sets from the training data, evaluates them on a validation set to select the best few-shot prompt. In contrast, the fine-tuning phase is formulated as a continuous optimization problem, where supervised fine-tuning is applied to update the model parameters for better task adaptation. Different from BetterTogether, we first transform the discrete–continuous optimization into a fully continuous problem, then propose a knowledge sharing mechanism and a supervised regularization loss to enable effective end-to-end co-optimization.
@@ Appendix C Additional Experiments
C.1 Case Study
@@
We present a case from the MATH dataset in Figure 8 to illustrate the advantages of MetaTuner over existing methods. Compared with pure prompt optimization approaches such as BPO, which can produce step-by-step reasoning but often generate overly lengthy intermediate steps and occasionally commit calculation errors that lead to incorrect answers, MetaTuner produces reasoning traces that are both accurate and concise while strictly adhering to the required output format. This improvement stems from its joint optimization of prompts and parameters, which allows the model to better adapt to downstream tasks. In contrast, when compared with traditional fine-tuning approaches such as DPO, MetaTuner also demonstrates clear advantages. Although DPO can arrive at the correct numerical solution, it fails to comply with formatting requirements and tends to generate relatively rigid and less context-aware reasoning. MetaTuner, by dynamically generating high-quality prompts that activate relevant internal knowledge while simultaneously refining model parameters, effectively combines the strengths of both prompt optimization and fine-tuning. As a result, MetaTuner achieves more stable, flexible, and diverse reasoning patterns, demonstrating both correctness and expressiveness.
@@
C.2 Efficiency Analysis
@@
During the inference stage, for a given query $x$, the MetaTuner framework needs to generate a set of prompts and parameters to complete the downstream task. Unlike the standard text generation task, there are additional costs associated with the LoRA parameter generation. In practice, we generate parameters for the o_proj part of the decoder layers in the downstream actor model $\mathcal{M}$. As mentioned in Section 3.4, each LoRA parameter is generated by a hyper-network $\phi_{q}=\{W_{d}^{b}\in\mathbb{R}^{d_{\mathcal{M}}\times l},W_{u}^{b}\in\mathbb{R}^{h_{\mathcal{G}}\times r_{\mathcal{M}}},W_{d}^{a}\in\mathbb{R}^{r_{\mathcal{M}}\times l},W_{u}^{a}\in\mathbb{R}^{h_{\mathcal{G}}\times k_{\mathcal{M}}}\}$. Therefore, depending on the task scenario and model choice, the additional cost varies with sequence length $l$, the hidden size $h_{\mathcal{G}}$ of the prompt generator, the LoRA rank $r_{\mathcal{M}}$ of the generated parameters, and the number of decoder layers $K$ in the downstream actor model. However, compared to the parameter quantity of the LLM itself, the hyper-network is relatively small, so it does not incur significant computational overhead. In Table 3, we present the inference costs on a single NVIDIA A100 80GB GPU, measuring the average time cost per query for parameter generation across different datasets and base models.
@@
Table 3: Average inference cost of parameter generation per query (second).
Base Model MATH GSM8K HotpotQA CosmosQA Avg
Qwen2.5-7B 0.13 0.11 0.02 0.30 0.14
Qwen2.5-3B 0.49 0.53 0.03 0.12 0.29
@@ D.1 Preliminary Experiment
We present the prompts used in the right subfigure of the preliminary experiment below. For constructing the training data, the prompt $p$ is concatenated in front of the query $x$ to form the model input, while the ground truth $y$ serves as the expected output. The supervised fine-tuning process is conducted using Llama-Factory (ref:llamafactory, ), with the epoch set to 1, the batch size set to 64, and the learning rate tuned over the values {5ee-4, 1ee-4}.
@@
D.2 Implementation of MetaTuner
Data splition. We split each dataset into training, validation, and test sets for model training and evaluation. For MATH, we use the 12K instances in the training set as specified in (ref:math_data_split, ), and randomly divide the remaining 500 instances into 50 for validation and 450 for testing. For GSM8K, the official test set contains 1,319 instances, so we allocate 473 instances for validation and the remaining 7K instances for training. For HotpotQA, we sample 20,000 instances from the official training set and randomly select 500 for validation and 2,000 for testing from the official dev-set-distractor. For CosmosQA, we use the entire 25,262 instances in the training set and randomly sample 500 and 2,000 instances from the test set for validation and testing, respectively.
Details of the warm-up stage. For the warm-up of $\mathcal{M}$, we use the queries $x$ from the original dataset as model inputs and the corresponding ground truth $y$ as model outputs for supervised fine-tuning. For the warm-up of $\mathcal{G}$, since there is no ground truth prompt, we construct the supervised training data with rejection sampling. Specifically, we first utilize GPT-4o to generate expert prompts $\mathcal{D}_{\text{expert}}=\{\hat{p}_{i}\}_{i=1}^{N}$, and then use the task reward $\mathcal{R}$ to filter out positive examples. The positive samples are then collected as $\mathcal{D}_{\text{PO}}=\{(x_{i},\hat{p}_{i})|\mathcal{R}(\mathcal{M}^{*}(\hat{p}_{i},x_{i}))=1\}_{i=1}^{N}$ for the supervised fine-tuning of prompt generation. Here, $\mathcal{M}^{*}$ and $\mathcal{G}^{*}$ represent the model after warm-up. Naturally, the pre-warmed model can form a pipeline $\hat{y_{i}}=\mathcal{M}^{*}(\mathcal{G}^{*}(\tilde{p},x_{i}),x_{i})$ to complete downstream tasks. We report the results in Table 4.
@@ D.2 Implementation of MetaTuner
Construction of the meta prompts. In this paper, since we adopt a prompt-rewrite paradigm for prompt generation, each meta prompt contains a query and an initial prompt, as well as an instruction to describe the prompt rewrite task. We follow the guidance of (ref:llm_dual_phase_po, ) to construct a base version of the prompt for each task as the initial prompt, which only includes a description of the task and the required answer format to facilitate answer parsing. For GSM8K, since the base model already produces highly consistent answer formats, we omit the answer format requirement in the initial prompt. The meta prompts on four datasets are presented below.

### v2

@@ 5 Related Work
Fine-tuning. Fine-tuning is a standard approach for adapting pre-trained language models to specific tasks. Recent advances in parameter-efficient fine-tuning (PEFT) scale this to LLMs by training small plug-in modules under a frozen backbone. Representative methods include Adapter Tuning (Houlsby et al., 2019), which inserts lightweight adapter layers, and LoRA (Hu et al., 2022), which applies low-rank updates to weight matrices. Building on these, recent works propose various post-training strategies (Ethayarajh et al., 2024; Guo et al., 2024; Jung et al., 2024; Xu et al., 2024; Yang et al., 2024a). For instance, RLHF (Li et al., 2023b) incorporates human feedback to improve alignment, while DPO (Rafailov et al., 2023) streamlines preference modeling through contrastive learning.
Prompt optimization. Prompt optimization has become a research hotspot with the rise of in-context learning. Early works focus on continuous prompt tuning (Li and Liang, 2021; Lester et al., 2022; Liu et al., 2024b; 2022; Lester et al., 2021), where learnable vectors are attached to the input. To mitigate the interpretability challenge (Wen et al., 2023), recent studies turn to discrete prompt optimization, which generates prompts in natural language. Existing approaches fall into two categories. The first employs LLMs as prompt optimizers (Liu et al., 2025b), typically in a feedback-driven manner, via strategies such as reflection-based refinement (Yang et al., 2024a; b; Austin and Chartock, 2024) or Monte Carlo Tree Search (MCTS) (Wang et al., 2024b; Pryzant et al., 2023; Liu et al., 2025a). The second (Kong et al., 2024; Juneja et al., 2024; Choi et al., 2024; Wang et al., 2024b; Chen et al., 2023) trains an auxiliary model to generate better prompts, most often with reinforcement learning, since prompts lack explicit ground truth and can be naturally guided by the task rewards.
Joint optimization. Recently, BetterTogether (Soylu et al., 2024) proposes to update prompts and model weights together to improve the task performance. In BetterTogether, prompt optimization and fine-tuning correspond to discrete and continuous optimization problems, respectively. During the prompt optimization phase, it samples candidate demonstration sets from the training data, evaluates them on a validation set to select the best few-shot prompt. In contrast, the fine-tuning phase is formulated as a continuous optimization problem, where supervised fine-tuning is applied to update the model parameters for better task adaptation. Different from BetterTogether, we first transform the discrete–continuous optimization into a fully continuous problem, then propose a knowledge sharing mechanism and a supervised regularization loss to enable effective end-to-end co-optimization.
@@ 6 Conclusion
Acknowledgement
This work is supported in part by National Natural Science Foundation of China (No. 62472427 and No. 62422215), Major Innovation & Planning Inter-disciplinary Platform for the “DoubleFirst Class” Initiative, Renmin University of China, Public Computing Cloud, Renmin University of China, fund for building world-class universities (disciplines) of Renmin University of China.
@@ Appendix C Additional Experiments
C.1 Main Results with More Datasets and Base Models
To further validate the effectiveness of MetaTuner across broader scenarios and model architectures, we conduct additional experiments on the Checkmate in One Move dataset using multiple base models, including Llama3.2-3B (Llama-3.2-3B-Instruct), Llama3.1-8B (Llama-3.1-8B-Instruct), and Mistral-7B (Mistral-7B-Instruct). The results in Table 3 show that MetaTuner consistently outperforms baseline methods, demonstrating the efficacy of our proposed prompt–parameter co-optimization framework, which is built upon parameter sharing and a supervised regularization loss.
Table 3: Main results with more datasets and base models.
Dataset Checkmate in One Move MATH
Base Model Qwen2.5-3B Qwen2.5-7B Llama3.2-3B Llama3.1-8B Llama3.2-3B Mistral-7B
Vanilla Methods
Base Model 0.00 0.00 0.00 21.33 25.11 10.89
Prompt Optimization Methods
RLPrompt 0.00 0.00 0.00 24.00 25.56 12.00
BPO 0.00 0.86 1.43 26.22 26.22 12.44
OPRO 0.00 0.57 1.14 25.77 26.22 11.78
CFPO 0.00 0.00 1.43 25.33 27.11 11.56
Fine-Tuning Methods
SFT 13.71 18.00 18.85 24.44 33.11 14.00
PPO 14.00 18.29 19.43 23.78 32.22 14.67
DPO 13.42 18.86 19.14 26.44 34.00 16.00
KTO 14.57 18.57 20.00 29.78 33.56 13.78
Hybrid Methods
BetterTogether 15.14 19.71 20.57 30.44 34.44 16.44
MetaTuner-J 17.71 20.86 21.43 31.78 18.44 35.78
C.2 Additional Generalization Analysis
We provide additional OOD experiments in Figure 9(a). The results show that MetaTuner achieves better out-of-distribution generalization compared to baseline methods. Theoretically, the generalization capability of our framework stems from the representation sharing mechanism. In MetaTuner, we can regard prompt generation and parameter generation as two distinct tasks and enforce them to share a common meta encoder. This shared-private multi-head architecture inherently induces a regularization effect (Park et al., 2019; Wang et al., 2022): the shared encoder must simultaneously support the learning objectives of multiple heads, compelling it to capture stable, core structural features that are invariant across tasks and independent of the specific training distribution. This aligns with classical theoretical results in multi-task learning (MTL) (Baxter, 2000; Maurer et al., 2016; D’Eramo et al., 2024; Bairaktari et al., 2023; Zhang et al., 2025): When multiple tasks share intermediate representations, the effective hypothesis space of the model is constrained, leading to tighter generalization error bounds. Intuitively, the model cannot learn highly task-specific, high-capacity representations independently for each task; instead, all tasks must find compatible representations within the same shared feature space. This capacity constraint significantly reduces overfitting to the idiosyncrasies of the training data distribution, thereby enhancing generalization across tasks and datasets. We believe that existing theoretical works on multi-task learning provide a solid foundation for understanding the generalization properties of MetaTuner.
(a) Additional generalization results. We select one dataset from the four (MATH, GSM8K, HotpotQA, CosmosQA) as the test set and use the remaining three as the training set.
C.3 Case Study
@@
Joint optimization for superior reasoning. We present a case from the MATH dataset in Figure 9 to illustrate the advantages of MetaTuner over existing methods. Compared with pure prompt optimization approaches such as BPO, which can produce step-by-step reasoning but often generate overly lengthy intermediate steps and occasionally commit calculation errors that lead to incorrect answers, MetaTuner produces reasoning traces that are both accurate and concise while strictly adhering to the required output format. This improvement stems from its joint optimization of prompts and parameters, which allows the model to better adapt to downstream tasks. In contrast, when compared with traditional fine-tuning approaches such as DPO, MetaTuner also demonstrates clear advantages. Although DPO can arrive at the correct numerical solution, it fails to comply with formatting requirements and tends to generate relatively rigid and less context-aware reasoning. MetaTuner, by dynamically generating high-quality prompts that activate relevant internal knowledge while simultaneously refining model parameters, effectively combines the strengths of both prompt optimization and fine-tuning. As a result, MetaTuner achieves more stable, flexible, and diverse reasoning patterns, demonstrating both correctness and expressiveness.
[…]
Learned prompts across different tasks. We provide several cases in Table LABEL:tab:learned_prompt_task to demonstrate the learned prompts across different tasks. From the examples, MetaTuner’s learned prompts exhibit clear cross-task commonalities and transferable patterns. In MATH and GSM8K, both learned prompts refine the initial CoT instruction by explicitly guiding step decomposition, intermediate computation, and clear explanation—despite differing problem types. In HotpotQA, the prompt preserves the core principle of concise, evidence-based answering, and sharpened to extract only task-relevant details while avoiding redundancy. In CosmosQA, the prompt emphasizes grounding the answer in contextual key details. In summary, despite differences in task-specific requirements, the prompts learned by MetaTuner consistently promote clear, logically structured responses and adapt their instructional details according to the nature of each task. This indicates that MetaTuner is able to identify and leverage shared structural patterns across tasks, while tailoring guidance to individual task needs, thereby enhances the transferability and adaptability of its prompts across diverse settings.
@@
Table 4: Cases of learned prompts with different base models.
Initial Prompt
[…]
Case 1
Query The lengths of two opposite sides of a square are decreased by 40% while the lengths of the other two sides are increased by 50% to form a rectangle. By what percent does the square’s area decrease?
[…]
Case 2
Query Find the greatest integer less than ($\sqrt{7}+\sqrt{5})^{6}$. (Do not use a calculator!)
[…]
Table 4: Cases of learned prompts with different base models. (Continued)
Table 5: Cases of learned prompts with Qwen2.5-7B across different tasks.
MATH
Query The lengths of two opposite sides of a square are decreased by 40% while the lengths of the other two sides are increased by 50% to form a rectangle. By what percent does the square’s area decrease?
[…]
GSM8K
Query Ted the T-Rex was planning to bring potato salad to the dinosaur picnic. He knows that an adult dinosaur will eat 10 lbs of potato salad, and a child will eat half as much as an adult. If there will be 20 adults and 5 children at the picnic, how many pounds of potato salad does Ted need to bring to the picnic if he hopes to have enough to feed everyone?
Initial Prompt Given a complex mathematical problem, solve it step by step using a chain-of-thought approach.
Learned Prompt Solve the following complex mathematical problem step by step using a chain-of-thought approach. Break down the problem into smaller parts, calculate intermediate values, and provide a clear explanation for each step.
CosmosQA
Query The black cloud’s clinic is full of complex patients who need three referrals apiece. The white cloud’s clinic is full of patients who don’t show up. What finally got me to believe in the power of the black cloud was one resident I worked with last year. One month she was working in the clinic but covering call in the PICU every fourth night (as I am doing this month). Why are you cover call in the PICU this month? A. The schedule for that job rotates monthly and new people have the responsibility of doing it monthly B. I am the best at covering calls at the PICU so I have that job C. None of the above choices. D. I needed to make extra money so I applied for the position
Initial Prompt Based on the provided context, select the most relevant and accurate option as the correct answer to the question. The answer should be directly related to the context, and must focus on the key details or events described.
Learned Prompt Based on the provided context, select the most relevant and accurate option as the correct answer to the question. Consider the details about the rotation of responsibilities and the nature of the job in the PICU. The answer should be directly related to the context, focusing on the key details or events described.
HotpotQA
Query (1) UKF is a brand owned by Luke Hood and AEI Media that focuses on sharing all aspects of electronic music. UKF Music was created by Luke Hood, who began sharing bass music through his original two YouTube Channels, UKF Drum & Bass and UKF Dubstep in 2009. On 28 June 2012, UKF celebrated one billion YouTube views since its creation. The UKF brand now has five YouTube Channels: UKF Music, UKF Dubstep, UKF Drum & Bass, UKF Mixes, and UKF Live. UKF has expanded beyond its YouTube Channels to creating compilation series and podcasts, organizing events, offering merchandise as well as its own ticketing platform. (2) Ysabella Brave is an American YouTube personality, artist, vocalist, singer and songwriter signed by Cordless Recordings, a division of the Warner Music Group. She was discovered through the popularity of her YouTube channels, Ysabella Brave and ysabellabravetalk. (3) Steven Jay Williams (born July 24, 1974), also known by his online alias Boogie2988 or just Boogie, is an American YouTube personality best known for his video rants about video games and nerd culture as a character named “Francis”. The Francis character is based on stereotypes of video game players and often parodies trending video game news, reaction and culture. Williams based the character on his early life experiences and has said that he wants viewers to hate the character for embodying gamer stereotypes. Boogie2988 videos range from absurd rants to serious discussions on daily life, such as the ethics of paid promotion on YouTube channels. He won the Trending Gamer award at The Game Awards 2016. (4) Joseph Michael Graceffa (born May 16, 1991) is an American YouTube personality, actor, author, producer, and singer. He has two active YouTube channels, both of which are named after him. One of them is for vlogging and the other is used for video gaming content. They both have a combined total of more than 1.7 billion views. His other YouTube channel, which he no longer uses, called WinterSpringPro, has 60 million views (as of August 2017). He was a contestant on the 22nd and 24th seasons of “The Amazing Race” and has appeared in short films (as web films) and short web television series on YouTube and its subscription-only service YouTube Red. Question: Shane and Friends have interviewed an American YouTube personality who has two active YouTube channels, which have a combined total of how many views?
Initial Prompt Use the relevant information provided to answer the multi-hop reasoning question. Please provide the most accurate and concise answer to the question based on the information provided. The answer should be directly from the dataset, with no extra details or explanations. For yes-no questions, respond with either ‘yes’ or ‘no.’ Avoid long or complex sentences.
Learned Prompt Use the relevant information provided to answer the multi-hop reasoning question. Identify the American YouTube personality mentioned and determine the combined total views of their two active YouTube channels. Provide the most accurate and concise answer directly from the dataset, without extra details or explanations. For yes-no questions, respond with either ‘yes’ or ‘no.’ Avoid long or complex sentences. Focus on extracting the specific information needed to answer the question accurately. Ensure the answer is clear and directly addresses the query.
Table 5: Cases of learned prompts with Qwen2.5-7B across different tasks. (Continued)
Table 6: Training cost of MetaTuner against prompt optimization and fine-tuning methods (hour).
Method MetaTuner PPO DPO BPO RLPrompt
Qwen2.5-7B $\approx$ 5 $\approx$3.5 $\approx$3 $\approx$3.5 $\approx$2.5
Qwen2.5-3B $\approx$3.5 $\approx$2.5 $\approx$2 $\approx$2 $\approx$1.5
C.4 Efficiency Analysis
Training efficiency. We take the MATH dataset as an example to compare the training overhead of MetaTuner’s joint training process against conventional prompt optimization and fine-tuning methods. All training procedures are carried out on two 80GB GPUs, and the associated costs are reported in Table 6. Since MetaTuner collects training samples online, to ensure a fair comparison, we include the time required for additional data collection in the training cost of the baselines as well. From the results, MetaTuner indeed incurs higher training costs compared to baseline approaches. However, since this training is performed only once, we consider the training cost to be practically acceptable.
@@ C.4 Efficiency Analysis
Inference efficiency. During the inference stage, for a given query $x$, MetaTuner needs to generate a set of prompts and parameters to complete the downstream task. Unlike the standard text generation task, there are additional costs associated with the LoRA parameter generation. In practice, we generate parameters for the o_proj part of the decoder layers in the downstream actor model $\mathcal{M}$. As mentioned in Section 3.4, each LoRA parameter is generated by a hyper-network $\phi_{q}=\{W_{d}^{b}\in\mathbb{R}^{d_{\mathcal{M}}\times l},W_{u}^{b}\in\mathbb{R}^{h_{\mathcal{G}}\times r_{\mathcal{M}}},W_{d}^{a}\in\mathbb{R}^{r_{\mathcal{M}}\times l},W_{u}^{a}\in\mathbb{R}^{h_{\mathcal{G}}\times k_{\mathcal{M}}}\}$. Therefore, depending on the task scenario and model choice, the additional cost varies with sequence length $l$, the hidden size $h_{\mathcal{G}}$ of the prompt generator, the LoRA rank $r_{\mathcal{M}}$ of the generated parameters, and the number of decoder layers $K$ in the downstream actor model. However, compared to the parameter quantity of the LLM itself, the hyper-network is relatively small, so it does not incur significant computational overhead.
In Table 7 and Table 8, we report the cost of generating prompts and LoRA parameters, respectively. For reference, we also present the original inference cost without generating prompts and LoRA parameters, as well as the average best performance for each category of methods. From the above results, we can see: (1) Compared to methods with prompt generation, MetaTuner achieves an absolute performance gain of 5.73% at the cost of an average additional inference latency of 0.22 seconds. (2) Compared to Vanilla and Fine-tuning methods, MetaTuner achieves an absolute performance gain of 10.51% at the cost of an average additional inference latency of 0.43 seconds.
@@
Table 7: Average inference cost of Qwen2.5-3B base model per query (second).
Dataset MATH GSM8K HotpotQA CosmosQA Avg Cost Avg Best Performance
Answer Generation 0.43 0.39 0.07 0.06 0.24 52.38
+Prompt 0.6(+0.17) 0.55(+0.16) 0.23(+0.16) 0.18(+0.12) 0.39(+0.15) 57.57
+Prompt & Parameter 0.73(+0.30) 0.66(+0.27) 0.25(+0.18) 0.48(+0.42) 0.53(+0.29) 65.10
Table 8: Average inference cost of Qwen2.5-7B base model per query (second).
Dataset MATH GSM8K HotpotQA CosmosQA Avg Cost Avg Best Performance
Answer Generation 0.71 0.68 0.14 0.11 0.41 60.30
+Prompt 0.99(+0.28) 0.95(+0.27) 0.44(+0.30) 0.36(+0.25) 0.69(+0.28) 64.68
+Prompt & Parameter 1.48(+0.77) 1.48(+0.80) 0.47(+0.33) 0.48(+0.37) 0.98(+0.57) 68.60
Improving inference efficiency. To accelerate the inference process, we further propose MetaTuner-C, which directly clusters the queries in the test set and then uses the prompt and parameters associated with the cluster centroid for all samples within that cluster during inference. Assuming the test set contains N samples and is partitioned into K clusters, the additional time overhead of MetaTuner-C compared to vanilla methods can be reduced to $t=K(t_{p}+t_{q})/N$, where $t_{p}$ and $t_{q}$ denote the time costs for generating the prompt and parameters for a single query, respectively. Notably, our main method, MetaTuner-J, can be viewed as a special case of MetaTuner-C when $K=N$. We evaluate MetaTuner-C with various $K$ using the Qwen2.5-3B base model on the MATH dataset and report the results in Table 9 and Table 10. Our analysis shows that MetaTuner-C achieves strong performance while significantly reducing inference time. For example, under the Qwen2.5-3B setting, using the best fine-tuning baseline DPO as a reference, MetaTuner-C with K=50 attains an absolute performance gain of 3.41% with only 0.03 seconds of additional latency per query.
Table 9: Performance of MetaTuner-C with Qwen2.5-3B model on MATH.
Method MetaTuner-J MetaTuner-C (K=50) MetaTuner-C (K=30) MetaTuner-C (K=10) DPO
Performance 41.33 40.00 39.11 37.56 36.59
Inference Cost (s) 0.73(+0.30) 0.46(+0.03) 0.45(+0.02) 0.437(+0.007) 0.43
Table 10: Performance of MetaTuner-C with Qwen2.5-7B model on MATH.
Method MetaTuner-J MetaTuner-C (K=50) MetaTuner-C (K=30) MetaTuner-C (K=10) DPO
Performance 48.67 45.56 45.11 44.89 43.78
Inference Cost (s) 1.48(+0.77) 0.80(+0.09) 0.76(+0.05) 0.73(+0.02) 0.71
@@ D.1 Preliminary Experiment
We present the prompts used in the right subfigure of the preliminary experiment below. For constructing the training data, the prompt $p$ is concatenated in front of the query $x$ to form the model input, while the ground truth $y$ serves as the expected output. The supervised fine-tuning process is conducted using Llama-Factory (Zheng et al., 2024), with the epoch set to 1, the batch size set to 64, and the learning rate tuned over the values {5ee-4, 1ee-4}.
@@
D.2 Implementation of MetaTuner
Data splition. We split each dataset into training, validation, and test sets for model training and evaluation. For MATH, we use the 12K instances in the training set as specified in (Lightman et al., 2024), and randomly divide the remaining 500 instances into 50 for validation and 450 for testing. For GSM8K, the official test set contains 1,319 instances, so we allocate 473 instances for validation and the remaining 7K instances for training. For HotpotQA, we sample 20,000 instances from the official training set and randomly select 500 for validation and 2,000 for testing from the official dev-set-distractor. For CosmosQA, we use the entire 25,262 instances in the training set and randomly sample 500 and 2,000 instances from the test set for validation and testing, respectively.
Details of the warm-up stage. For the warm-up of $\mathcal{M}$, we use the queries $x$ from the original dataset as model inputs and the corresponding ground truth $y$ as model outputs for supervised fine-tuning. For the warm-up of $\mathcal{G}$, since there is no ground truth prompt, we construct the supervised training data with rejection sampling. Specifically, we first utilize a leading closed-source LLM to generate expert prompts $\mathcal{D}_{\text{expert}}=\{\hat{p}_{i}\}_{i=1}^{N}$, and then use the task reward $\mathcal{R}$ to filter out positive examples. The positive samples are then collected as $\mathcal{D}_{\text{PO}}=\{(x_{i},\hat{p}_{i})|\mathcal{R}(\mathcal{M}^{*}(\hat{p}_{i},x_{i}))=1\}_{i=1}^{N}$ for the supervised fine-tuning of prompt generation. Here, $\mathcal{M}^{*}$ and $\mathcal{G}^{*}$ represent the model after warm-up. Naturally, the pre-warmed model can form a pipeline $\hat{y_{i}}=\mathcal{M}^{*}(\mathcal{G}^{*}(\tilde{p},x_{i}),x_{i})$ to complete downstream tasks. We report the results in Table 11.
To further investigate the impact of generated expert data during the warm-up stage, we additionally study how using a smaller expert dataset (in Table 12) or a weaker expert model (in Table 13) affects the MetaTuner’s final performance. Our analysis reveals that MetaTuner exhibits robust performance under both weaker and smaller expert dataset settings. As the size of the expert dataset decreases, performance remains stable both after warm-up and joint training. When a weaker expert model is used for warm-up, the post-warm-up performance degrades moderately as the expert model’s capability diminishes. However, this performance gap is largely mitigated after joint training. This suggests that the warm-up phase only requires a basic level of prompt-generation competence. The proposed MetaTuner framework can subsequently discover superior prompt–parameter combinations during joint training phase, thereby enhancing overall performance.
Table 12: Performance of Qwen2.5-3B on MATH with smaller expert datasets.
Data Quantity 100% 60% 20%
Performance After Warm-up 35.33 36.67 34.44
Performance After Joint Training 41.33 41.33 40.89
Table 13: Performance of Qwen2.5-3B on MATH with weaker expert datasets.
Expert Model Proprietary LLM Qwen2.5-32B Qwen2.5-14B
Performance After Warm-up 35.33 33.11 31.33
Performance After Joint Training 41.33 40.22 40.00
Construction of the meta prompts. In this paper, since we adopt a prompt-rewrite paradigm for prompt generation, each meta prompt contains a query and an initial prompt, as well as an instruction to describe the prompt rewrite task. We follow the guidance of (Yang et al., 2024b) to construct a base version of the prompt for each task as the initial prompt, which only includes a description of the task and the required answer format to facilitate answer parsing. For GSM8K, since the base model already produces highly consistent answer formats, we omit the answer format requirement in the initial prompt. The meta prompts on four datasets are presented below.